OSWG: Documents Catalog

Deb Richardson

deb@oswg.org

This catalog is an organized listing of the documents currently available through the Open Source Writers Group. Updates are made as often as possible.

1. Copyright and Licensing

Copyright (c) 1999 by Deb Richardson. This material may be distributed only subject to the terms and conditions set forth in the Open Publication License v0.4 (8 June 1999) or later. The current version of the License is available at http://www.opencontent.org/openpub/.

2. Catalog organization and feedback

This catalog is organized according to the subject matter of the documents, then listed alphabetically according to title. Later versions of this document may contain additional organizational structures.

If you notice any information in this catalog that is inaccurate or out of date, please send a note to <deb@oswg.org>.

Contributions, additions, or updates are always welcome, and can also be sent to

<deb@oswg.org>.

3. Documentation and documentation tools

3.1. DocBook, A Practical Introduction

[HTML (single page) (http://www.oswg.org/oswgnightly/oswg/en_US.ISO_8859-1/articles/DocBook-Intro/docbook-intro.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/DocBook-Intro/docbook-intro/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/docbook-intro.html)]

Author: Deb Richardson <deb@oswg.org>

Version: 0.1 incomplete

Last Revised: November 14, 1999

A practical introduction to using the DocBook SGML DTD for marking up technical (and other) documentation, written with the new user in mind.

4. Linux developer documentation

4.1. GCC HOWTO

[HTML (single page) (http://www.oswg.org/oswg-nightly/oswg/en_GB.ISO_8859-1/books/linux-c-programming/GCC-HOWTO.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_GB.ISO_8859-1/books/linux-c-programming/GCC-HOWTO/t1.html)] [Other formats (http://www.oswg.org/oswg-nightly/GCC-HOWTO.html)] Author: Dan Barlow <dan@telent.net>

Version: n/a

Last Revised: May, 1999

This document covers how to set up the GNU C compiler and development libraries under Linux, and gives an overview of compiling, linking, running and debugging programs under it.

4.2. PA-RISC/Linux Boot HOWTO

[HTML (single page) (http://www.oswg.org/oswgnightly/oswg/en_US.ISO_8859-1/articles/parisc-linux/parisc-linux-boot.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/parisc-linux/parisc-linux-boot/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/parisc-linux-boot.html)]

Author: Deb Richardson <deb@thepuffingroup.com>

Version: 0.3 (draft)

Last Revised: November 25, 1999

This document outlines the procedure for getting the current PA-RISC/Linux development kernel to boot on your PA-RISC system. This HOWTO does not currently include the information needed for booting your system from disk. The booting from disk procedures will be added as soon as they are available.

4.3. RPM HOWTO, aka: RPM at Idle

[HTML (single page) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/RPM-HOWTO/RPM-HOWTO.html)] [HTML (multiple pages)

(http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/RPM-HOWTO/RPM-HOWTO/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/RPM-HOWTO.html)]

Author: Donnie Barnes <djb@redhat.com>

Version: 3.0

Last Revised: November, 1999

RPM is the "RPM Package Manager". It is an open packaging system available for anyone to use. This document describes how to use RPM to manage software packages on your system, as well as how to use RPM to package your software for use with RPM.

4.4. SCSI Drivers, An Introduction to

[HTML (single page) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/alan-cox/scsi-drivers/scsi-drivers.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/alan-cox/scsi-drivers/scsi-drivers/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/scsi-drivers.html)]

Author: Alan Cox <alan@redhat.com>

Version: n/a

Last Revised: 1999

Writing a driver for a simple SCSI controller under Linux. The Linux kernel does most of the work for SCSI devices, so a "dumb" SCSI interface can usually be relatively painless to write.

5. Linux user documentation

5.1. Booting the Kernel

[HTML (single page) (http://www.oswg.org/oswgnightly/oswg/en_US.ISO_8859-1/articles/alessandro-rubini/boot/boot.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/alessandro-rubini/boot/boot/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/boot.html)]

Author: Alessandro Rubini <rubini@linux.it>

Version: n/a

Last Revised: 1999

This article describes the steps that are performed to boot the Linux kernel. While this kind of information is not relevant to the system’s functionality, it’s interesting to see how the different architectures bring the system up.

5.2. Cable Modem Providers HOWTO

[HTML (single page) (http://www.oswg.org/oswgnightly/oswg/en_US.ISO_8859-1/articles/Cable-Modem/Cable-Modem.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/Cable-Modem/Cable-Modem/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/Cable-Modem.html)]

Author: Vladimir Vuksan <vuksan@veus.hr>

Version: 4.8

Last Revised: 24 November, 1999 This document attempts to answer basic questions on how to connect your Linux box to a cable modem or cable Internet provider.

5.3. DHCP mini-HOWTO

[HTML (single page) (http://www.oswg.org/oswgnightly/oswg/en_US.ISO_8859-1/articles/DHCP/DHCP.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/DHCP/DHCP/t1.html)] [Other formats (http://www.oswg.org/oswg-nightly/DHCP.html)]

Authors: Vladimir Vuksan <vuksan@veus.hr>

Version: 4.6

Last Revised: 17 November, 1999

This document attempts to answer basic questions on how to set up your Linux box to serve as a DHCP server or a DHCP client.

5.4. Exchange Server Replacement HOWTO

[HTML (single page) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/exchange-replacement-howto/exchange-replacement-howto.html)] [HTML (multiple pages) (http://www.oswg.org/oswgnightly/oswg/en_US.ISO_8859-1/articles/exchange-replacement-howto/exchange-replacement-howto/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/exchange-replacement-howto.html)]

Author: Curt Johnson <cjohnson@arrayservices.com>

Version: n/a

Last Revised: January, 2000 This document describes the installation and configuration of an IMAP and POP3 mail server using LDAP as the user database.

5.5. Gnome Display Manager Reference Manual

[HTML (single page) (http://www.oswg.org/oswgnightly/oswg/en_US.ISO_8859-1/articles/gdm-reference/gdm-reference.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/gdm-reference/gdm-reference/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/gdm-reference.html)]

Author: Martin K. Petersen <mkp@mkp.net>

Version: n/a

Last Revised: November, 1999

GDM is a replacement for XDM, the X Display Manager. Unlike its competitors (X3DM, KDM, WDM) GDM was written from scratch and does not contain any original XDM / X Consortium code.

5.6. Lilo mini-HOWTO

[HTML (single page) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/Lilo-Mini-Howto/lilo-mini-howto.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/Lilo-Mini-Howto/lilo-mini-howto/t1.html)] [Other formats (http://www.oswg.org/oswg-nightly/lilo-mini-howto.html)]

Authors: Cameron Spitzer <cls@truffula.sj.ca.us>, Alessandro Rubini

<rubini@linux.it>

Version: 2.03

Last Revised: 7 November, 1999

LILO is the most used Linux Loader for the x86 flavour of Linux. This mini-HOWTO describes some typical LILO installations, and is intended as a supplement to the LILO User’s Guide.

5.7. LVS Cluster Configuration HOWTO

[HTML (single page) (http://www.oswg.org/oswgnightly/oswg/en_US.ISO_8859-1/articles/cluster-howto/cluster-howto.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/cluster-howto/cluster-howto/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/cluster-howto.html)]

Author: Wayne Sherrill <sherrill@redhat.com>

Version: n/a

Last Revised: 18 November, 1999

This document explains how to set up and administer a Linux Virtual Server (LVS) cluster that provides highly available Web and FTP services.

5.8. Online Troubleshooting Resources: HOWTO

[HTML (single page) (http://www.oswg.org/oswgnightly/oswg/en_US.ISO_8859-1/articles/tronlineDOC/tronlineDOC.html)] [HTML (multiple pages) (http://www.oswg.org/oswgnightly/oswg/en_US.ISO_8859-1/articles/tronlineDOC/tronlineDOC/t1.html)] [Other formats (http://www.oswg.org/oswg-nightly/tronlineDOC.html)]

Author: Howard Mann <howardm@xmission.com>

Version: 1.1

Last Revised: 8 November, 1999

This document directs Linux users to resources available on the Internet that provide access to a vast amount of Linux-related information useful in troubleshooting problems.

5.9. PHHTTPD

[HTML (single page) (http://www.oswg.org/oswgnightly/oswg/en_US.ISO_8859-1/articles/phhttpd/phhttpd.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/phhttpd/phhttpd/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/phhttpd.html)]

Author: Zach Brown <zab@zabbo.net>

Version: n/a

Last Revised: January, 2000

phhttpd is an HTTP acclerator. It serves fast static HTTP fetches from a local file-system and passes slower dynamic requests back to a waiting server. It features a lean networking I/O core and an aggressive content cache that helps it perform its job efficiently.

5.10. RPM HOWTO, aka: RPM at Idle

[HTML (single page) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/RPM-HOWTO/RPM-HOWTO.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/RPM-HOWTO/RPM-HOWTO/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/RPM-HOWTO.html)] Author: Donnie Barnes <djb@redhat.com>

Version: 3.0

Last Revised: November, 1999

RPM is the "RPM Package Manager". It is an open packaging system available for anyone to use. This document describes how to use RPM to manage software packages on your system, as well as how to use RPM to package your software for use with RPM.

5.11. Visual Bell HOWTO

[HTML (single page) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/alessandro-rubini/visual-bell/visual-bell-howto.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/alessandro-rubini/visual-bell/visual-bell-howto/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/visual-bell-howto.html)]

Author: Alessandro Rubini <rubini@linux.it>

Version: 2.2

Last Revised: November, 1999

This document explains how to use termcap to configure a visual bell on one’s system and describes how to disable audible bells on demand.

6. Mailing list software and related documentation

6.1. Advice for Mailing List Owners

[HTML (single page) (http://www.oswg.org/oswgnightly/oswg/en_US.ISO_8859-1/articles/mladvice/mladvice.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/mladvice/mladvice/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/mladvice.html)]

Author: Kirrily Robert <skud@netizen.com.au>

"This is a collection of advice for mailing list owners based on my experience of running technical and social lists of all sizes since 1994. It’s not a technical document, and you won’t find out how to configure your mailing list software here. Instead, you’ll learn some general techniques to help you make your mailing list run smoothly and keep yourself sane in the process."

7. Open Source Writers Group: project documentation

7.1. OSWG: Document Catalog

[HTML (single page) (http://www.oswg.org/oswgnightly/oswg/en_US.ISO_8859-1/articles/OSWG-Catalog/oswg-catalog.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/OSWG-Catalog/oswg-catalog/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/oswg-catalog.html)]

Author: Deb Richardson <deb@oswg.org>

Version: 1.0

Last Revised: November 15, 1999

An organized listing of the documents currently available through the Open Source Writers Group. Updated as often as possible and/or required.

7.2. OSWG: Licensing Policy

[HTML (single page) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/OSWG-Licensing-Policy/oswg-licensing-policy.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/OSWG-Licensing-Policy/oswg-licensing-policy/t1.html)] [Other formats (http://www.oswg.org/oswg-nightly/oswg-licensing-policy.html)]

Author: Deb Richardson <deb@oswg.org>

Version: 0.1

Last Revised: November, 1999

This document has been written in order to simplify the process of deciding which licenses are allowable for documents that are to be included as part of the OSWG documentation set.

7.3. OSWG: Mission Statement

[HTML (single page) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/OSWG-Mission-Statement/oswg-mission-stmt.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/OSWG-Mission-Statement/oswg-mission-stmt/t1.html)] [Other formats (http://www.oswg.org/oswg-nightly/oswg-mission-stmt.html)]

Author: Deb Richardson <deb@oswg.org>

Version: 0.1

Last Revised: November, 1999

The OSWG Mission Statement has been written to clarify the purpose, goals, and future of the OSWG. This document exists to help other projects understand what the OSWG is about, and to help current and future OSWG members understand what we’re trying to do.

7.4. OSWG: Roadmap

[HTML (single page) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/OSWG-Roadmap/oswg-roadmap.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/OSWG-Roadmap/oswg-roadmap/t1.html)] [Other formats (http://www.oswg.org/oswg-nightly/oswg-roadmap.html)]

Author: Deb Richardson <deb@oswg.org>

Version: 0.1

Last Revised: November, 1999

This document outlines the major short-term and long-term goals and plans of the Open Source Writers Group.

8. Open Source: Philosophy, Theory, Practice, and History

8.1. Cathedrals, Bazaars, and the Town Council

[HTML (single page) (http://www.oswg.org/oswgnightly/oswg/en_US.ISO_8859-1/articles/alan-cox/town-council/tcouncil.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/alan-cox/town-council/tcouncil/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/tcouncil.html)]

Author: Alan Cox <alan@redhat.com>

Year Published: 1999

"These are some of my thoughts on the Bazaar model that I figure are worth sharing. It’s also a guide to how to completely screw up a free software project."

8.2. I don’t care if space aliens ate my mouse (Porting Linux to the Apple Macintosh)

[HTML (single page) (http://www.oswg.org/oswgnightly/oswg/en_US.ISO_8859-1/articles/alan-cox/mac/mac.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/alan-cox/mac/mac/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/mac.html)]

Author: Alan Cox <alan@redhat.com>

Year Published: 1999

"A study in both the technical and human issues in porting the Linux OS to a new M68K target platform. And an analysis of the effectiveness of the platform independant features of the Macintosh. Mostly however an essay on the dangers of saying "Well all you need to do is..."

8.3. The Risks of Closed Source Computing

[HTML (single page) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/alan-cox/risks/risks-closed-source.html)] [HTML (multiple pages)

(http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/alan-cox/risks/risks-closed-source/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/risks-closed-source.html)]

Author: Alan Cox <alan@redhat.com>

Year Published: 1999

One large vendor likes to talk about the risks of Open Source software, but the strange thing is, the risk is actually in closed technologies. This article looks at the real risks in following a proprietary software path.

8.4. Software Libre and Commercial Viability

[HTML (single page) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/alessandro-rubini/viability/viability.html)] [HTML (multiple pages) (http://www.oswg.org/oswg-nightly/oswg/en_US.ISO_8859-1/articles/alessandro-rubini/viability/viability/index.html)] [Other formats (http://www.oswg.org/oswg-nightly/viability.html)]

Author: Alessandro Rubini <rubini@linux.it>

Year Published: 1998

An article that addresses some of the social and commercial implications of free software.